I am stumped, i know you plug in the bottom formula but trying to solve it has me confused
g(x)=x^2+3x

h(x)=(x+3)/2

find (h*g)(-1)



Answer by stanbon(75887)

---
g(x)=x^2+3x

h(x)=(x+3)/2

find (h*g)(-1)
============================
Multiply the two algebraic expressions; then evaluate that product at x=-1.
--------------------------------------------
(h*g)(-1) = [(x+3)/2](-1)
---
[(x^3+3x^2)/2](-1)
---
[(-1)^3+3(-1)^2]/2
---
=[-1+3]/2
=[2/2]
= 1
==============
Cheers,
Stan H.
=============================